OpenStudy (anonymous):

(3)/(x+2)+(4x+5)/(3x+6)

OpenStudy (thomas5267):

Is this the question? \[ \frac{3}{x+2}+\frac{4x+5}{3x+6} \]

OpenStudy (anonymous):

yes

OpenStudy (thomas5267):

\[ \begin{align*} &\dfrac{3}{x+2} + \dfrac{4x+5}{3x+6} \\ =&\dfrac{3}{x+2} +\dfrac{4x+5}{3(x+2)} &\text{Factor as much as possible.} \\ =&\dfrac{3}{(x+2)} \times \dfrac{3}{3}+\dfrac{4x+5}{3(x+2)} &\dfrac{3}{3}=1 \\ =&\dfrac{9}{3x+2}+\dfrac{4x+2}{3(x+2)} \end{align*} \] Finish the last step on your own.

OpenStudy (anonymous):

where did you get the 3/3?

OpenStudy (thomas5267):

I have to times it by \(\frac{3}{3}\) in order to add the first term and second term together. Remember that you have to get the denominator the same before adding fractions up?

OpenStudy (anonymous):

oh okay